OpenStudy (anonymous):

Apply the translations y = f(x + 5) - 4 to f(x) = (x - 2)^2

OpenStudy (anonymous):

Are you looking for like a ratio?

OpenStudy (anonymous):

That answer would be Impossible.

OpenStudy (anonymous):

fx+f5-4 to f(x)=x-4

OpenStudy (anonymous):

I am going to contact my teacher, thank you for trying.

OpenStudy (anonymous):

There's probably something wrong with my assignment booklet where I took this question from. I sent my teacher an e-mail and she will contact me to either help me solve it or inform me if it is an error.

hartnn (hartnn):

your next maths questions should go in 'Mathematics' section, there you'll get help quickly. to get f(x+5) from f(x), just replace 'x' in f(x) by 'x+5' so, f(x+5) = (x+5-2)^2 = (x+3)^2 then y= f(x+5)-4 = (x+3)^2 -4 if you want you can simplify that to get the form y= ax^2+bx+c